User name Armstrong
Log entry time 11:45:36 on September22,2005
Entry number 155599
keyword=left septum trip.
The Left septum tripped. Why? No beam; perhaps while Ed is
cooling right Q1 the additional cryogen draw needed stole some
coolant from the septum. The coli temperatures are 11.5 K, but seems
cooling reasonably rapidly.
Disable audio alarms while we wait to reset the septum.
I guess if a septum has to trip, having it happen when we are not
taking any data is a good time...